\begin{thebibliography}{99}
\bibitem{atanassov1} K.T. Atanassove, Intuitionistic Fuzzy Sets,  Springer Physica-Verlag.Heidelberg (1999).
\bibitem{atanassov}  K.T. Atanassov, Intuitionistic fuzzy sets, Fuzzy Sets and Systems 	20(1986) 87-96.
	
\bibitem{babitha}  K.V. Babitha, S.J. Johan,  Hesitant fuzzy soft sets,  J. New Results Sci.		3(2013) 98-107.
	
\bibitem{borah} M.J. Borah, B. Hazarika, Some operators on interval-valued Hesitant Fuzzy soft sets, arXiv:1604.00902v1, 1 Apr 2016.
	
	
\bibitem{broumi} S. Broumi and F. Smarandache, New operations over interval valued intuitionistic hesitant fuzzy set, Math. Stat. 2(2) (2014) 62-71.

\bibitem{cagman}  N. \c{C}a\v{g}man and N. S. Engino\v{g}lu, Soft Set Theory
and Uni-int Decision Making, Euro. J. Opera. Research  207(2010) 848-855.
\bibitem{chang}  C. L. Chang, Fuzzy Topological Spaces, J. Math. Anal. Appl.
	24(1)(1968) 182-190.
\bibitem{chen}  N.Chen, Z.Xu, M.Xia,  Interval-valued hesitant preference relations and their applications to group decision making, Knowledge-Based Systems 37(2013) 528-540.


\bibitem{Chiao} Kuo-Ping Chiao, Fundamental properties of interval vector
max-norm, Tamsui Oxford J. Math. 18(2)(2002) 219-233.

\bibitem{Dwyer51} P.S.Dwyer, Linear Computation, New York, Wiley, 1951.

%\bibitem{Dwyer53} P.S.Dwyer, Errors of matrix computation, simultaneous equations and eigenvalues, National Bureu of Standarts, Applied Mathematics Series, 29(1953), 49-58.

%\bibitem{Fischer} P.S.Fischer, Automatic propagated and round-off error analysis, paper presented at the 13th National Meeting of the Association of Computing Machinary, June 1958.

\bibitem{liao} H.C. Liao, Z.S. Xu, Subtraction and divison operations over hesitant fuzzy sets,  J. Intell. Fuzzy Systems 27(1)(2014) 65-72.
\bibitem{mahanta} J. Mahanta and P. K. Das, Results on Fuzzy soft topological Spaces,  arXiv:1203.0634v1. % {\bf math.GM}(3 Mar 2012).

\bibitem{majietal1} P. K. Maji, R. Biswas, R. Roy, An application of soft sets in a decision making
problem,  Comput. Math. Appl.  44(2002) 1077-1083.

\bibitem{majietal2} P. K. Maji, R. Biswas, R. Roy, Soft set theory, Comput. Math. Appl. 45(2003) 555-562.

\bibitem{maji}  P. K. Maji, R. Biswas and A. R. Roy, Fuzzy Soft
 Sets,  J. Fuzzy Math. 9(3)(2001) 589-602.

\bibitem{molodstov}  D. A. Molodstov, Soft Set Theory-First Result,  Comput. Math. Appl.
 37(1999) 19-31.

\bibitem{MolodtsovLeonov} D. A. Molodtsov, V. Y. Leonov, D.V. Kovkov,  Soft sets technique and its application,  Nechetkie Sistemy i Myagkie Vychisleniya  1(1)(2006) 8-39.


%\bibitem{Moore} R.E.Moore, Automatic Error Analysis in Digital Computation, LSMD-48421, Lockheed Missiles and Space Company, 1959.

%\bibitem{MooreYang62} R.E.Moore and C.T.Yang, Interval Analysis I, LMSD-285875, Lockheed Missiles and Space Company, 1962.

%\bibitem{MooreYang58} R.E.Moore and C.T.Yang,\ Theory of an interval algebra and its application to numeric analysis, RAAG Memories II, Gaukutsu Bunken Fukeyu-kai, Tokyo, 1958.
\bibitem{neog} T. J. Neog, D. K. Sut and G. C. Hazarika, Fuzzy Soft Topological Spaces, Inter. J. Latest Trends Math. 	2(1)(2012) 54-67.
	
\bibitem{pengyang} X. Peng, Y. Yang, Interval-valued Hesitant Fuzzy Soft Sets and their Application in Decision Making,  Fundamenta Informaticae 141(1)(2015) 71-93.
	
	
\bibitem{ray}  S. Ray and T. K. Samanta, A note on Fuzzy Soft Topological Spaces,  Annals  Fuzzy Math. Inform. 	3(2)(2012) 305-311.
\bibitem{shabir}  M. Shabir and M. Naz, On soft topological spaces,  Comput. Math. Appl.
	61(2012) 412-418.
\bibitem{tanay}  B. Tanay and M. Bur\c{c} Kandemir, Topological Structure of fuzzy soft sets,  Comput. Math. Appl.
	61(2011) 2952-2957.
\bibitem{torra}  V. Torra, Hesitant fuzzy sets, Int. J. Intell. Systems
25(6)(2010) 529-539.

 \bibitem{torra2} V. Torra and Y. Narukawa, On hesitant fuzzy sets and decision.
{\it In proceeding of the 18th IEEE internatiuonal conference on fuzzy systems, Jeju Island, Republic of Korea} (2009) 1378-1382.


\bibitem{verma}  R. Verma and B.D.Sharma, New operations over hesitant fuzzy sets,  Fuzzy Inform. Eng. 2(2013) 129-146.


\bibitem{wang} J. Wang, X. Li and X. Chen,  Hesitant fuzzy soft sets with application in multicrteria group decision making problems,  The Scientific World J. 2015(2015) 1-14.


\bibitem{xia} M. Xia, Z. Xu, Hesitant fuzzy information aggregation in decision making, Int. J. Appr. Reasoning 52(2011) 395-407.

\bibitem{xu91} Z. Xu, Intuitionistic fuzzy aggregation operators,  IEEE Transaction on Fuzzy Systems  15(6)(2007) 1179-1187.

\bibitem{xu9} Z. Xu, Q.L. Da,  The uncertain OWA operator,  Int. J. Intell. Systems 17 (2002) 569-575.


\bibitem{xu} Z. Xu, M. Xia,  Distance and similrity measures for hesitant fuzzy sets,  Inform. Sci. 181 (2011) 2128-2138.


\bibitem{zadeh}  L. A. Zadeh, Fuzzy Sets,  Inform. Control. 8(1965) 338-353.

\bibitem{zhang}  H. Zhang, L. Xiong and W. Ma, On interval-Valued Hesitant Fuzzy Soft Sets,  Math. Problems  Eng. 2015(2015), Article ID 254764, 17 pages. %doi.org/10.1155/2015/254764



\end{thebibliography}

\documentclass[10pt,twoside]{amsart}
\usepackage{amsmath, amsthm, amscd, amsfonts, amssymb, graphicx}
\newtheorem{thm}{Theorem}[section]

\newtheorem{prop}[thm]{Proposition}
\newtheorem{defn}[thm]{Definition}

\newtheorem{problem}[thm]{Example}
\numberwithin{equation}{section}

\begin{document}


\begin{center}\large{{\bf{Notes on interval-valued Hesitant fuzzy soft Topological Space
				}}}

\vspace{0.5cm}

{\bf{
Manash Jyoti Borah$^{1}$ and Bipan Hazarika$^{2\ast}$
}}

\vspace{.2cm}
$^{1}$Department of Mathematics, Bahona College,  Jorhat-785 101, Assam, India\\ Email:mjyotibora9@gmail.com

\vspace{.2cm}

$^{2}$Department of Mathematics, Rajiv Gandhi University, Rono Hills, Doimukh-791 112, Arunachal Pradesh, India\\ Email:  bh\_rgu$@$yahoo.co.in
\end{center}
\vspace{.5cm}
\title{}
\author{}
\thanks{\today\\{$^{\ast}$The corresponding author}}

\begin{abstract}
 In this paper we introduce the notion of interval valued hesitant fuzzy soft topological space. Also the concepts of interval valued hesitant fuzzy soft closure, interior and neighbourhood are introduced here and established some important results.  \\

 Keywords: Fuzzy soft sets, Interval-valued Hesitant fuzzy sets, Interval-valued Hesitant fuzzy soft sets, Interval-valued Hesitant fuzzy soft topological space.

AMS subject classification no: 03E72.
\end{abstract}

\maketitle
\pagestyle{myheadings}
\markboth{\rightline {\scriptsize Borah, Hazarika}}
         {\leftline{\scriptsize Notes on interval-valued Hesitant fuzzy...}}

\maketitle

\section{Introduction}
The concept of interval arithmetic was first suggested by Dwyer \cite{Dwyer51} in
1951. 
 Chiao in \cite{Chiao} introduced sequence of interval numbers and
defined usual convergence of sequences of interval number.  A set consisting of a closed interval of real numbers x such that $a\leq x\leq
b$ is called an interval number. A real interval can also be considered as a
set. Thus we can investigate some properties of interval numbers, for instance
arithmetic properties or analysis properties. We denote the set of all real
valued closed intervals by I$%
\mathbf{R}
.$ Any elements of I$%
\mathbf{R}
$ is called closed interval and denoted by $\overline{x}.$ That is
$\overline{x}=\left\{  x\in%
\mathbb{R}
:\text{ }a\leq x\leq b\right\}.$ An interval number $\overline{x}$ is a
closed subset of real numbers (see \cite{Chiao}). Let $x_{l}$ and $x_{r}$ be
first and last points of $\overline{x}$ interval number, respectively. For
$\overline{x}_{1},\overline{x}_{2}\in$I$%
\mathbf{R}
,$ we have $\overline{x}_{1}=\overline{x}_{2}\Leftrightarrow x_{1_{l}}%
$=$x_{2_{l}},$ $x_{1_{r}}$=$x_{2_{r}}.$ $\overline{x}_{1}+\overline{x}%
_{2}=\left\{  x\in%
\mathbb{R}
:x_{1_{l}}+x_{2_{l}}\leq x\leq x_{1_{r}}+x_{2_{r}}\right\},$ and if
$\alpha\geq0,$ then $\alpha\overline{x}=\left\{  x\in%
\mathbb{R}
:\text{ }\alpha x_{1_{l}}\leq x\leq\alpha x_{1_{r}}\right\}  $ and if
$\alpha<0,$ then $\alpha\overline{x}=\left\{  x\in%
\mathbb{R}
:\text{ }\alpha x_{1_{r}}\leq x\leq\alpha x_{1_{l}}\right\}  ,$
\[
\overline{x}_{1}.\overline{x}_{2}=\left\{
\begin{array}
[c]{c}%
x\in%
\mathbb{R}
:\min\left\{  x_{1_{l}}.x_{2_{l}},x_{1_{l}}.x_{2_{r}},x_{1_{r}}.x_{2_{l}%
},x_{1_{r}}.x_{2_{r}}\right\}  \leq x\\
\leq\max\left\{  x_{1_{l}}.x_{2_{l}},x_{1_{l}}.x_{2_{r}},x_{1_{r}}.x_{2_{l}%
},x_{1_{r}}.x_{2_{r}}\right\}
\end{array}
\right\}  .
\]
The most appropriate theory for
dealing with uncertainties is the theory of fuzzy sets, introduced by L.A. Zadeh \cite{zadeh} in 1965.
This theory brought a paradigmatic change in mathematics. But there exists difficulty, how to
set the membership function in each particular case. 
The Hesitant fuzzy set, as one of the extensions of Zadeh \cite{zadeh} fuzzy set, allows the membership degree that an element to a set
presented by several possible values, and  it can express the hesitant information
more comprehensively than other extensions of fuzzy set. Torra and Narukawa \cite{torra2} introduced the
concept of hesitant fuzzy set. Xu and Xia \cite{xu}
defined the concept of hesitant fuzzy element, which can be considered as
the basic unit of a hesitant fuzzy set, and is a simple and effective tool used to express the
decision makers’ hesitant preferences in the process of decision making. So many  researchers  has done lots of research work on aggregation, distance,
similarity and correlation measures, clustering analysis, and decision making with
hesitant fuzzy information. Babitha and John \cite{babitha} defined another important soft set i.e. Hesitant fuzzy soft sets. They introduced basic operations such as intersection, union, compliment  and De Morgan's law was proved. Chen et al. \cite{chen} extended hesitant fuzzy sets into interval-valued hesitant fuzzy environment and introduced the concept of interval-valued hesitant fuzzy sets. Zhang et al. \cite{zhang} introduced some operations such as complement, "AND","OR", ring sum and ring product on interval-valued hesitant fuzzy soft sets.    \\

There are many theories like theory of probability, theory of fuzzy sets, theory of
intuitionistic fuzzy sets, theory of rough sets etc. which can be considered as mathematical
tools for dealing with uncertain data, obtained in various fields of engineering, physics,
computer science, economics, social science, medical science, and of many other diverse
fields. But all these theories have their own difficulties. The theory of intuitionistic fuzzy sets (see \cite{atanassov1, atanassov}) is a
more generalized concept than the theory of fuzzy sets, but this theory has the same
difficulties. All the above mentioned theories are successful to some extent in dealing with
problems arising due to vagueness present in the real world. But there are also cases where
these theories failed to give satisfactory results, possibly due to inadequacy of the
parameterization tool in them. As a necessary supplement to the existing mathematical tools
for handling uncertainty,  Molodtsov \cite{molodstov} introduced the theory of soft sets as
a new mathematical tool to deal with uncertainties while
modelling the problems in engineering, physics, computer
science, economics, social sciences, and medical sciences.
Molodtsov et al \cite{MolodtsovLeonov} successfully applied soft sets in
directions such as smoothness of functions, game theory,
operations research, Riemann integration, Perron integration, probability, and theory of measurement. Maji et al \cite{majietal1} gave the first practical application of soft sets in
decision-making problems.  Maji et al \cite{majietal2} defined
and studied several basic notions of the soft set theory. Also \c{C}a\v{g}man et al \cite{cagman} studied several basic notions of the soft set theory.  V. Torra \cite{torra,torra2} and Verma and Sharma \cite{verma} discussed the relationship between hesitant fuzzy set and showed that the envelope of hesitant fuzzy set is an intuitionistic fuzzy set. Zhang et al \cite{zhang} introduced weighted interval-valued hesitant fuzzy soft sets and finally applied it in decision making problem.\\
The notion of topological space is defined on crisp sets and hence
it is affected by different generalizations of crisp sets like
fuzzy sets and soft sets. In 1968,
C. L. Chang \cite{chang} introduced fuzzy topological space and in
2011, subsequently \c{C}a\v{g}man et al.  \cite{cagman} and Shabir et al.
\cite{shabir} introduced fuzzy soft topological spaces and
studied neighborhood.  Mahanta et al.   \cite{mahanta},
Neog et al. \cite{neog} and Ray et al.   \cite{ray} introduced fuzzy
soft topological spaces in different direction.\\
In this paper, in section 3, First we give a counter example of equality of IVHFSSs proposed by Zhang et al.	\cite{zhang}. Secondly we point out that proposition 3.11 in a previous paper by Borah and Hazarika \cite{borah} true in general by counter example. Thirdly we introduce about notion of topological space. 
\section{Preliminaries and Definitions}

In this section we recall some basic concepts and definitions
regarding fuzzy soft sets, hesitant fuzzy  set and hesitant fuzzy soft set.

\begin{defn}
\cite{maji} Let $U$ be an initial universe and $F$ be a set of parameters. Let $\tilde{P}(U)$ denote the power set of $U$ and $A$ be a
non-empty subset of $F.$  Then $F_A$ is called a
fuzzy soft set over $U,$ where $F:A\rightarrow \tilde{P}(U)$ is a
mapping from $A$ into $\tilde{P}(U).$
\end{defn}

\begin{defn}
\cite{molodstov}  $F_E$ is called a soft
set over $U$ if and only if $F$ is a mapping of $E$ into the set of all
subsets of the set $U.$
\end{defn}
In other words, the soft set is a parameterized family of subsets
of the set $U.$ Every set $F(\epsilon),$ $\epsilon\tilde{\in} E,$
from this family may be considered as the set of
$\epsilon$-element of the soft set $F_E$  or as the set of
$\epsilon$-approximate elements of the soft set.
\begin{defn}
	\cite{atanassov,xu91} Let intuitionistic fuzzy value IFV(X) denote the family of all IFVs defined on the universe X, and let $ \alpha, \beta\in IFV(X) $ be given as: \\
	$ \alpha=(\mu_\alpha,\nu_\alpha), \beta=(\mu_\beta,\nu_\beta),$
	\begin{enumerate}
\item[	(i)] $ \alpha\cap \beta=(\min(\mu_\alpha,\mu_\beta ), \max(\nu_\alpha,\nu_\beta )) $
\item[	(ii)] $ \alpha\cup \beta=(\max(\mu_\alpha,\mu_\beta ), \min(\nu_\alpha,\nu_\beta )) $	
	\item[(iii)] $ \alpha \ast \beta=(\frac{\mu_\alpha+\mu_\beta}{2(\mu_\alpha.\mu_\beta+1)},\frac{\nu_\alpha+\nu_\beta}{2(\nu_\alpha.\nu_\beta+1)} ).$	
	\end{enumerate}

\end{defn}
\begin{defn}
	\cite{torra} Given a fixed set $X, $ then a hesitant fuzzy set (shortly HFS) in $X$ is in terms of a function that when applied to $X$ return a subset of $[0, 1].$ We express the HFS by a mathematical symbol:\\
	$ F=\{<h, \mu_{F}(x)> : h\in X\}, $
	where $ \mu_{F}(x) $ is a set of some values in $[0,1],$ denoting the possible membership degrees of the element $ h\in X $ to the set $ F.$ $ \mu_{F}(x) $ is called a hesitant fuzzy element (HFE) and $ H $ is the set of all HFEs. 
\end{defn}
\begin{defn}
	\cite{torra} Let $ \mu_{1},  \mu_{2}\in H $ and three operations are defined as follows:
	\begin{enumerate}
\item[(1)] $ \mu_{1}^C=\cup_{\gamma_{1}\in\mu_{1}} \{1-\gamma_{1}\}; $
	
	\item[(2)] $ \mu_{1}\cup\mu_{2}=\cup_{\gamma_{1}\in\mu_{1}, \gamma_{2}\in\mu_{2}}\max \{\gamma_{1},\gamma_{2} \} ; $
	
	\item[(3)] $ \mu_{1}\cap\mu_{2}=\cap_{\gamma_{1}\in\mu_{1}, \gamma_{2}\in\mu_{2}}\min \{\gamma_{1},\gamma_{2} \}.$
			\end{enumerate}
\end{defn}

\begin{defn}
\cite{chen} Let $X$ be a reference set, and $D[0,1]$ be the set of all closed subintervals of $[0,1].$ An IVHFS on $X$ is 
	$ F=\{<h_i, \mu_{F}(h_i)> : h_i\in X, i=1,2,...n\}, $
	where $\mu_{F}(h_i):X\rightarrow D[0,1]$ denotes all possible interval-valued membership degrees of the element $ h_i\in X $ to the set F. For convenience, we call $ \mu_{F}(h_i) $ an interval-valued hesitant fuzzy element (IVHFE), which reads
	$\mu_{F}(h_i)=\{\gamma: \gamma\in\mu_{F}(h_i) \}.$\\
Here $ \gamma=[\gamma^L, \gamma^U] $ is an interval number. $ \gamma^L=\inf \gamma $ and $ \gamma^U=\sup \gamma $ represent the lower and upper limits of $ \gamma,$ respectively. An IVHFE is the basic unit of an IVHFS and it can be considered as a special case of the IVHFS. The relationship between IVHFE and IVHFS is similar to that between interval-valued fuzzy number and interval-valued fuzzy set. 
\end{defn}
\begin{problem}\label{exam36}
	Let $ U=\{h_1, h_2\} $ be a reference set and let 
	$ \mu_F(h_1)=\{[0.6,0.8],[0.2,0.7]\},\mu_F(h_2)=\{[0.1,0.4]\}  $ be the IVHFEs of $ h_i (i=1,2) $ to a set $F,$ respectively. Then IVHFS  $F$ can be written as  
	 $F=\{<h_1,\{[0.6, 0.8], [0.2, 0.7]\}>, <h_2, \{[0.1, 0.4]\}>\}.$
\end{problem}

\begin{defn}
	\cite{xu9} Let $ \tilde{a}=[\tilde{a}^L,\tilde{a}^U ] $ and 
$ \tilde{b}=[\tilde{b}^L,\tilde{b}^U ] $ be two interval numbers and $ \lambda\geq0, $ then 

\begin{enumerate}
	\item[(i)] $\tilde{a}=\tilde{b}\Leftrightarrow\tilde{a}^L=\tilde{b}^L$ and $ \tilde{a}^U=\tilde{b}^U ;$
	\item[(ii)]$\tilde{a}+\tilde{b}=[\tilde{a}^L+\tilde{b}^L, \tilde{a}^U+\tilde{b}^U] ;$
	\item[(iii)]$ \lambda\tilde{a}=[\lambda\tilde{a}^L,\lambda\tilde{a}^U ] ,$ especially $ \lambda\tilde{a}=0, $ if $ \lambda=0.$ 
\end{enumerate}
\end{defn}

\begin{defn}\label{defn209}
	\cite{xu9} Let $ \tilde{a}=[\tilde{a}^L,\tilde{a}^U ] $ and 
	$ \tilde{b}=[\tilde{b}^L,\tilde{b}^U ], $ and let $ l_a=\tilde{a}^U-\tilde{a}^L $ and $ l_b=\tilde{b}^U-\tilde{b}^L; $ then the degree of possibility of $\tilde{a}\geq\tilde{b}  $ is formulated by
	\[p(\tilde{a}\geq\tilde{b})=\max\left\{1-\max\left(\frac{\tilde{b}^U-\tilde{a}^L}{l_{\tilde{a}}+l_{\tilde{b}}},0\right),0\right\}.\]
	Above equation is proposed in order to compare two interval numbers, and to rank all the input arguments.
\end{defn}
\begin{defn}\label{defn210}
	\cite{chen} For an IVHFE $ \tilde{\mu} ,$ $ s(\tilde{\mu}) =\frac{1}{l_{ \tilde{\mu}}}\sum_{ \tilde{\gamma}\in \tilde{\mu}}\tilde{\gamma}$ is called the score function of $ \tilde{\mu} $ with $l_{ \tilde{\mu}}$ being the number of the interval values in  $ \tilde{\mu}, $ and $ s(\tilde{\mu}) $ is an interval value belonging to $[0,1].$ For two IVHFEs $\tilde{\mu_1}  $ and $\tilde{\mu_2} , $ if $  s(\tilde{\mu_1})\geq s(\tilde{\mu_2}),$ then $\tilde{\mu_1}\geq \tilde{\mu_2}. $\\
We can judge the magnitude of two IVHFEs using above equation. 
\end{defn}
\begin{defn}
\cite{chen} Let $\tilde{\mu}, \tilde{\mu_1}$ and $ \tilde{\mu_2}$ be three IVHFEs, then 
\begin{enumerate}
	\item[(i)]$\tilde{\mu}^C=\{[1-\tilde{\gamma}^U,1-\tilde{\gamma}^L]:\tilde{\gamma}\in\tilde{\mu} \} ; $
\item[(ii)] $\tilde{\mu_1}\cup\tilde{\mu_2}=\{[\max(\tilde{\gamma_1}^L,\tilde{\gamma_2}^L),\max(\tilde{\gamma_1}^U,\tilde{\gamma_2}^U)]:\tilde{\gamma_1}\in\tilde{\mu_1}, \tilde{\gamma_2}\in\tilde{\mu_2} \} ; $
\item[(iii] $\tilde{\mu_1}\cap\tilde{\mu_2}=\{[\min(\tilde{\gamma_1}^L,\tilde{\gamma_2}^L),\min(\tilde{\gamma_1}^U,\tilde{\gamma_2}^U)]:\tilde{\gamma_1}\in\tilde{\mu_1}, \tilde{\gamma_2}\in\tilde{\mu_2} \} ; $
\item[(iv)] $\tilde{\mu_1}\oplus \tilde{\mu_2} =\{[\tilde{\gamma_1}^L+\tilde{\gamma_2}^L-\tilde{\gamma_1}^L.\tilde{\gamma_2}^L, \tilde{\gamma_1}^U+\tilde{\gamma_2}^U-\tilde{\gamma_1}^U.\tilde{\gamma_2}^U]:\tilde{\gamma_1}\in\tilde{\mu_1}, \tilde{\gamma_2}\in\tilde{\mu_2} \} ;  $
\item[(v)] $\tilde{\mu_1}\otimes \tilde{\mu_2} =\{[\tilde{\gamma_1}^L.\tilde{\gamma_2}^L,\tilde{\gamma_1}^U.\tilde{\gamma_2}^U]:\tilde{\gamma_1}\in\tilde{\mu_1}, \tilde{\gamma_2}\in\tilde{\mu_2} \}.$  
\end{enumerate}
\end{defn}
\begin{prop}
\cite{chen} For three IVHFEs $ \tilde{\mu}, \tilde{\mu_1} $ and $ \tilde{\mu_2} ,$ we have 
\begin{enumerate}
	\item[(i)]$\tilde{\mu_1}^C\cup\tilde{\mu_2}^C= (\tilde{\mu_1}\cap\tilde{\mu_2})^C; $
	
	\item[(ii)]$ \tilde{\mu_1}^C\cap\tilde{\mu_2}^C= (\tilde{\mu_1}\cup\tilde{\mu_2})^C; $
\end{enumerate}
\end{prop}
\begin{defn}
	\cite{wang} Let $U$ be an initial universe and $E$ be a set of parameters. Let $\tilde{F}(U)$ be the set of all hesitant fuzzy subsets of $ U. $ Then $F_E$ is called a hesitant fuzzy soft set (HFSS) over $ U, $ where $\tilde {F}:E\rightarrow \tilde{F}(U). $\\
	A HFSS is a parameterized family of hesitant fuzzy subsets of $U,$ that is, $\tilde{F}(U).$ For all $\epsilon\tilde{\in} E,$  $F(\epsilon)$ is referred to as the set of $ \epsilon-$approximate elements of the HFSS $F_E.$ It can be written as 	
	\[\tilde{F}(\epsilon)=\{<h, \mu_{\tilde{F}(\epsilon)(x)}> : h\in U\}.\]
	 Since  HFE can represent the situation, in which different membership function are considered possible (see \cite{torra}), $ \mu_{\tilde{F}(\epsilon)(x)} $ is a set of several possible values, which is the hesitant fuzzy membership degree. In particular, if $ \tilde{F}(\epsilon) $ has only one element,  $ \tilde{F}(\epsilon)$  can be called a hesitant fuzzy soft number. For convenience, a hesitant fuzzy soft number (HFSN) is denoted by $ \{<h, \mu_{\tilde{F}(\epsilon)(x)}>\}.$ 
	 
\end{defn}
\begin{problem}\label{exam36}
		Suppose $ U=\{h_1,h_2\} $ be an initial universe and $ E=\{e_1, e_2, e_3, e_4\} $ be a set of parameters. Let $ A=\{e_1, e_2\} .$ Then the hesitant fuzzy soft set $ F_A $ is given as \\ $F_A=\{F(e_1)=\{<h_1,\{0.6, 0.8\}>, <h_2, \{0.8, 0.4, 0.9\}>\},  F(e_2)=\{<h_1,\{0.9, 0.1, 0.5\}>, <h_2,\{0.2 \}>\}\}.$
	
\end{problem}

\begin{defn}
\cite{zhang} Let $ (U,E) $ be a soft universe and $ A\subseteq E .$ Then $ F_A $ is called an interval valued hesitant fuzzy soft set over $U,$ where $ F $ is a mapping given by $F:A \rightarrow IVHF(U). $\\
An interval-valued hesitant fuzzy soft set is a parameterized family of interval-valued hesitant fuzzy subset of $ U. $ That is to say, $ F(e) $ is an interval-valued hesitant fuzzy subset in $ U,\forall e\in A. $ Following the standard notations, $ F(e) $ can be written as 
  \[\tilde{F}(e)=\{<h, \mu_{\tilde{F}(e)(x)}> : h\in U\}.\]
\end{defn}

\begin{problem}\label{exam36}
		Suppose $ U=\{h_1,h_2\} $ be an initial universe and $ E=\{e_1, e_2, e_3, e_4\} $ be a set of parameters. Let $ A=\{e_1, e_2\} .$ Then the interval valued hesitant fuzzy  soft set $ F_A $ is given as
	
 $F_A=\{e_1=\{<h_1,[0.6, 0.8]>, <h_2, [0.1, 0.4]>\}\\ 
  e_2=\{<h_1,[0.2, 0.6],[0.3,0.9] >, <h_2, [0.2,0.5],[0.2,0.8], [0.2,0.8]>\}\}.$
\end{problem}

\begin{defn}
\cite{zhang} Let $ U $ be an initial universe and let $ E $ be a set of parameters. Supposing that $ A, B \tilde{\subseteq}E, F_A $and $ F_B $ are two interval-valued hesitant fuzzy soft sets, one says that $ F_A $ is an interval-valued hesitant fuzzy soft subset of $ G_B $ if and only if 
\begin{enumerate}
\item[(i)] $ A\tilde{\subseteq}B, $
\item[(ii)]$\gamma_1^{\sigma(k)} \tilde{\leq}\gamma_2^{\sigma(k)},$
\end{enumerate}
where for all $ e\tilde{\in}A,~ x\tilde{\in}U,~ \gamma_1^{\sigma(k)} $and $\gamma_2^{\sigma(k)}  $ stand for the kth largest interval number in the IVHFEs $~ \mu_{F(e)(x)} $ and $\mu_{ G(e)(x)}, $ respectively. In this case, we write $ F_A\tilde{\subseteq}G_A. $
\end{defn}
\begin{defn}
	\cite{zhang} The complement of $ F_A, $ denoted by $ F_A^C, $ is defined by \[F_A^C(e)=\{<h, \mu_{\tilde{F}^C(e)(x)}> : h\in U\},\]  where $ \mu_{F}^C:A\rightarrow IVHF(U) $ is a mapping given by $\mu_{\tilde{F}^C(e)}, \forall e\tilde{\in}A $ such that $\mu_{\tilde{F}^C(e)} $ is the complement of interval-valued hesitant fuzzy element $ \mu_{\tilde{F}(e)}  $ on $ U. $
\end{defn}
\begin{defn}
	\cite{zhang} An interval-valued hesitant fuzzy soft set is said to be an empty interval-valued hesitant fuzzy soft set, denoted by $ \tilde{\phi},$ if $ F:E\rightarrow IVHF(U)$ such that \[\tilde{F}(e)=\{<h, \mu_{\tilde{F}(e)(x)}> : h\in U\}=\{<h, \{[0,0]\}> : h\in U\}, \forall e\tilde{\in}E. \]
\end{defn}

\begin{defn}
	\cite{zhang} An interval-valued hesitant fuzzy soft set is said to be an full interval-valued hesitant fuzzy soft set, denoted by $ \tilde{E},$ if $ F:E\rightarrow IVHF(U)$ such that 
\[\tilde{F}(e)=\{<h, \mu_{\tilde{F}(e)(x)}> : h\in U\}=\{<h, \{[1,1]\}> : h\in U\}, \forall e\tilde{\in}E. \]
\end{defn}
\begin{defn}
	\cite{borah}
		The union of two interval-valued hesitant fuzzy soft sets $ F_A $ and $ G_B $ over $ (U,E) ,$ is the interval-valued hesitant fuzzy soft set $ H_C ,$ where $ C=A\cup B $ and,  $\forall e\tilde{\in}C,$

		$\mu_{H(e)}=\begin{cases}
		\mu_{F(e)}, & \textrm{if $e\tilde{\in}A-B;$}\\
		\mu_{G(e)}, & \textrm{if $ e\tilde{\in}A-B;$}\\
		\mu_{F(e)}\cup \mu_{G(e)}, & \textrm{if $e\tilde{\in}A\cap B.$}\end{cases}$
		
		We write $ F_A \tilde{\cup}G_B=H_C.$
\end{defn}
\begin{defn}
	\cite{borah}  The intersection of two interval-valued hesitant fuzzy soft sets $ F_A $ and $ G_B $ with $ A\cap B\neq\phi $ over $ (U,E) ,$ is the interval-valued hesitant fuzzy soft set $ H_C ,$ where $ C=A\cap B, $ and,  $\forall e\tilde{\in}C, \mu_{ H(e)}=\mu_{F(e)}\cap \mu_{G(e)}.$ 
	We write $ F_A \tilde{\cap}G_B=H_C.$	
\end{defn}
\section{Interval-valued hesitant fuzzy soft topological space}
In this section, First we give a counter example of equality of IVHFSSs proposed by Zhang et al.	\cite{zhang}. Secondly we point out that proposition 3.11 in a previous paper by Borah and Hazarika \cite{borah} true in general by counter example. Thirdly we introduce about notion of topological space. 
\begin{defn}
	\cite{zhang} Let $ F_A $ and $ G_B $ be two interval-valued hesitant fuzzy soft sets. Now $ F_A $ and $ G_B $ are said to be interval-valued hesitant fuzzy soft equal if and only if  
	\begin{enumerate}
		\item[(i)] $ F_A\tilde{\subseteq}G_B, $
		\item[(ii)]$ G_B\tilde{\subseteq}F_A, $
	\end{enumerate}
	which can be denoted by $ F_A=G_B $
\end{defn}
\begin{problem}
	$F_A=\{e_1=\{<h_1,[0.2, 0.5]>, <h_2, [0.5, 0.8],[0.4,0.9]>\}, \\ e_2=\{<h_1,[0.3, 0.6], [0.4,0.8]>, <h_2, [0.6,0.8]>\}\}.$\\
	
	$G_A=\{e_1=\{<h_1,[0.2, 0.5],[0.2,0.5]>, <h_2, [0.5, 0.8],[0.4,0.9]>\},\\  e_2=\{<h_1,[0.3, 0.6], [0.4,0.8]>, <h_2, [0.6,0.8],[0.6,0.8]>\}\}.$\\
	Therefore  $  F_A\tilde{\subseteq}G_A $ and $ G_A\tilde{\subseteq}F_A. $\\
	Hence $ F_A=G_A $
\end{problem}
\begin{prop}\label{prop38}
	Let 	$ F_A, G_B$ and $ H_C $ be three interval-valued hesitant fuzzy soft sets. Then the following  are satisfied:
	
	\begin{enumerate}
		\item[(i)] $F_A\tilde{\cup}(G_B\tilde{\cap}H_C)=(F_A\tilde{\cup} G_B)\tilde{\cap}(F_A\tilde{\cup}H_C)$
		\item[(ii)]$F_A\tilde{\cap}(G_B\tilde{\cup}H_C)=(F_A\tilde{\cap} G_B)\tilde{\cup}(F_A\tilde{\cap}H_C).$
		
	\end{enumerate}
\end{prop}

\begin{proof}
	We consider IVHFSSs.\\
		$F_A=\{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.8],[0.5,0.6],[0.3,0.6]>\},\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0] >, <h_2, [0.8,1.0],[0.2,0.6]>\}\}.$\\
		    $G_B=\{e_1=\{<h_1,[0.7, 0.9],[0.0,0.6]>, <h_2, [0.4, 0.7],[0.4,0.5]>\}, \\ e_2=\{<h_1,[0.6, 0.8]>, <h_2, [0.3,0.8],[0.3,0.6]>\},\\
		e_3=\{<h_1,[0.5, 0.6],[0.3,0.6] >, <h_2, [0.1,0.6],[0.3,0.9],[0.3,0.6]>\}\}.$\\
	and\\
	  $H_C=\{e_2=\{<h_1,[0.4, 0.6],[o.2,0.6],[0.7,1.0]>, <h_2,[0.3,0.8],>\},\\  e_3=\{<h_1,[0.2, 0.5],[0.3,0.5] >, <h_2, [0.6,0.8],[0.2,0.5]>\}\}.$\\

	(i)  We have\\
	$ F_A \tilde{\cup}G_B= \{e_1=\{<h_1,[0.3, 0.8],[0.7,0.9]>, <h_2, [0.4, 0.6],[0.4,0.8][0.5,0.7]>\}\\  e_2=\{<h_1,[0.6, 0.9][0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\}\\
		e_3=\{<h_1,[0.3, 0.6],[0.5,0.6] >, <h_2, [0.1,0.6],[0.3,0.9],[0.3,0.6]>\}\}.$\\	
		$F_A\tilde{\cup}H_C= \{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8][0.5,0.6]>\}\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0],[0.7,1.0]>, <h_2, [0.3,0.8],[0.8,1.0]>\}\\
	e_3=\{<h_1,[0.2, 0.5],[0.3,0.5] >, <h_2, [0.2,0.5],[0.6,0.8]>\}\}.$\\
	
	$(F_A\tilde{\cup} G_B)\tilde{\cap}(F_A\tilde{\cup}H_C)\\
	=\{e_1=\{<h_1,[0.3, 0.8],[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8][0.5,0.6]>\}\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0],[0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\}\\
	e_3=\{<h_1,[0.2, 0.5],[0.3,0.5] >, <h_2, [0.1,0.5],[0.3,0.8],[0.3, 0.6]>\}\}.$\\
	Again\\
	$G_B\tilde{\cap}H_C= \{e_2=\{<h_1,[0.2, 0.6],[0.4,0.6],[0.6,0.8]>, <h_2, [0.3,0.6],[0.3,0.8]>\}\\
	e_3=\{<h_1,[0.2, 0.5],[0.3,0.5] >, <h_2, [0.1,0.5],[0.3,0.8],[0.3,0.6]>\}\}.$\\
	Therefore\\ $F_A\tilde{\cup}(G_B\tilde{\cap}H_C)$\\
	$ =\{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8][0.5,0.6]>\}\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0],[0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\}\\
	e_3=\{<h_1,[0.2, 0.5],[0.3,0.5] >, <h_2, [0.1,0.5],[0.3,0.8],[0.3, 0.6]>\}\}.$\\
	Hence $F_A\tilde{\cup}(G_B\tilde{\cap}H_C) =(F_A\tilde{\cup} G_B)\tilde{\cap}(F_A\tilde{\cup}H_C).$\\
	
	(ii) We have \\
	$ F_A \tilde{\cap}G_B= \{e_1=\{<h_1,[0.0, 0.6],[0.3,0.8]>, <h_2, [0.3, 0.5],[0.3,0.7][0.4,0.6]>\}\\  e_2=\{<h_1,[0.2, 0.8][0.6,0.8]>, <h_2, [0.2,0.6],[0.3,0.8]>\}\}.$\\
	$G_B\tilde{\cup}H_C= \{e_1=\{<h_1,[0.0, 0.6],[0.7,0.9]>, <h_2, [0.4, 0.5],[0.4,0.7][0.4,0.7]>\}\\  e_2=\{<h_1,[0.6, 0.8],[0.6,0.8],[0.7,1.0]>, <h_2, [0.3,0.8],[0.3,0.8]>\}\\
	e_3=\{<h_1,[0.3, 0.6],[0.5,0.6] >, <h_2, [0.2,0.6],[0.6,0.9],[0.6,0.8]>\}\}.$\\
	Therefore\\ $F_A\tilde{\cap}(G_B\tilde{\cup}H_C)$\\
	$ =\{e_1=\{<h_1,[0.0, 0.6],[0.3,0.8]>, <h_2, [0.3, 0.5],[0.3,0.7][0.4,0.6]>\}\\  e_2=\{<h_1,[0.2, 0.8],[0.6,0.8],[0.7,1.0]>, <h_2, [0.2,0.6],[0.3,0.8]>\}\}.$\\
	
	Again\\
	$F_A\tilde{\cap}H_C= \{e_2=\{<h_1,[0.2, 0.6],[0.4,0.6],[0.7,1.0]>, <h_2, [0.2,0.6],[0.3,0.8]>\}\}.$\\ 
	Therefore \\
	$(F_A\tilde{\cap} G_B)\tilde{\cup}(F_A\tilde{\cap}H_C)$\\
	$=\{e_1=\{<h_1,[0.0, 0.6],[0.3,0.8]>, <h_2, [0.3, 0.5],[0.3,0.7][0.4,0.6]>\}\\  e_2=\{<h_1,[0.2, 0.8],[0.6,0.8],[0.7,1.0]>, <h_2, [0.2,0.6],[0.3,0.8]>\}\}.$\\ 
	
	Hence $F_A\tilde{\cap}(G_B\tilde{\cup}H_C) =(F_A\tilde{\cap} G_B)\tilde{\cup}(F_A\tilde{\cap}H_C).$ 
\end{proof}
\begin{defn}  A interval-valued hesitant fuzzy soft topology $\tau$ on	$(U,E)$ is a family of interval-valued hesitant fuzzy soft sets over $(U,E)$ satisfying the following properties:
	\begin{enumerate}
		\item[(i)] $\tilde{\phi},\tilde{E}\tilde{\in}\tau$
		\item[(ii)] if $F_{A},G_{B}\tilde{\in}\tau,$  then
		$F_{A}\tilde{\cap}G_{B}\tilde{\in}\tau.$
		\item[(iii)] if ${F_{A_{\alpha}}\tilde{\in}}\tau$ for all
		$\alpha\tilde{\in}\Delta$ an index set, then
		$\bigcup_{\alpha\in{\Delta}}F_{A_{\alpha}}\tilde{\in}\tau.$
	\end{enumerate}
\end{defn}
\begin{problem}\label{prob35} Let $ U=\{h_1,h_2\} $ and $ E=\{e_1, e_2, e_3\},$ and consider 
	$ A=\{e_1, e_2, e_3\},\\ B=\{e_1, e_2\}\tilde{\subseteq}E.$\\
	Let
	$F_A=\{e_1=\{<h_1,[0.7,0.9],[0.3, 0.8]>, <h_2, [0.4, 0.6],[0.5,0.7],[0.4,0.8]>\}\\  e_2=\{<h_1,[0.6, 0.9], [0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\}\\
	e_3=\{<h_1,[0.3, 0.6],[0.5,0.6] >, <h_2,[0.3,0.9], [0.1,0.6]>\}\}.$\\
		$G_B=\{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3,0.8],[0.3, 0.6]>\}\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0] >, <h_2,[0.8,1.0], [0.2,0.6]>\}\}.$\\
	Now rearrange the membership value of	$ F_A$ and $ G_B$  with the help of  Definitions \ref{defn209}, \ref{defn210} and assumptions given by \cite{chen}, we have\\
	$F_A=\{e_1=\{<h_1,[0.3, 0.8],[0.7,0.9]>, <h_2, [0.4, 0.6],[0.4,0.8],[0.5,0.7]>\}\\  e_2=\{<h_1,[0.6, 0.9], [0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\}\\
	e_3=\{<h_1,[0.3, 0.6],[0.5,0.6] >, <h_2, [0.1,0.6],[0.3,0.9]>\}\}.$\\
	
	$G_B=\{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8]>\}\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0] >, <h_2, [0.2,0.6],[0.8,1.0]>\}\}.$\\
	Suppose a collection $ \tau $ of interval-valued hesitant fuzzy soft sets over $ (U,E) $ as 
	$\tau=\{\tilde{\phi}, \tilde{E}, \tilde{F_A}, \tilde{G_B}\}.$\\
	 	Therefore \\
		(i) $ \tilde{\phi} , \tilde{E} \tilde{\in} \tau$\\
	(ii) $\tilde{\phi} \tilde{\cap} \tilde{E}=\tilde{\phi}, \tilde{\phi} \tilde{\cap} F_A=\tilde{\phi},\tilde{\phi} \tilde{\cap} G_B=\tilde{\phi}, \tilde{E} \tilde{\cap} F_A=F_A, \tilde{E} \tilde{\cap} G_B=G_B $ and \\
	$ F_A \tilde{\cap}G_B=\{e_1=\{<h_1,[0.3, 0.8],[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8], [0.3,0.8]>\}\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0] >, <h_2, [0.2,0.6],[0.8,1.0]>\}\}.$ \\
		Hence	$  F_A \tilde{\cap}G_B=G_B. $\\
		(iii) $\tilde{\phi} \tilde{\cup} \tilde{E}=\tilde{E}, \tilde{\phi} \tilde{\cup} F_A=F_A,\tilde{\phi} \tilde{\cup} G_B=G_B, \tilde{E} \tilde{\cup} F_A=\tilde{E}, \tilde{E} \tilde{\cup} G_B=\tilde{E} $ and \\
	$ F_A \tilde{\cup}G_B=\{e_1=\{<h_1,[0.3, 0.8],[0.7,0.9]>, <h_2, [0.4, 0.6],[0.4,0.8],[0.5,0.7]>\}\\  e_2=\{<h_1,[0.6, 0.9], [0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\}\\
	e_3=\{<h_1,[0.3, 0.6],[0.5,0.6] >, <h_2, [0.1,0.6],[0.3,0.9]>\}\}=F_A,$ \\ and
	$ \tilde{\phi} \tilde{\cup} \tilde{E}\tilde{\cup}F_A=\tilde{E}, \tilde{\phi} \tilde{\cup} \tilde{E}\tilde{\cup}G_B=\tilde{E}, \tilde{E}\tilde{\cup}F_A\tilde{\cup} G_B=\tilde{E},\tilde{\phi} \tilde{\cup}\tilde{E}\tilde{\cup}F_A\tilde{\cup} G_B=\tilde{E}  $\\
	Therefore $ \tau $ is a IVHFS topology on $(U,E).$
	
\end{problem}

\begin{defn} 
If $ \tau $ is a IVHFS topology on $ (U,E) ,$ the triple $ (U, E, \tau) $ is said to be a interval-valued hesitant fuzzy soft topological space (IVHFSTS). Also each member of $ \tau $ ia called a interval-valued hesitant fuzzy soft open set in $(U,E, \tau).  $ 
\end{defn}
\begin{problem} 
From example \ref{prob35}, The triple $ (U,E,\tau) $ is a IVHFS topological space  and the interval-valued hesitant fuzzy soft open sets in $ (U,E,\tau) $ are $\tilde{\phi}, \tilde{E}, \tilde{F_A}, \tilde{G_B}.$
\end{problem}
\begin{problem} 
A IVHFSS $F_A$ over $ (U,E) $ is called an interval-valued hesitant fuzzy soft closed set in $ (U,E,\tau) $ if and only if its complement  $F_A^C$ is a  interval-valued hesitant fuzzy soft open set in $ (U,E,\tau). $
\end{problem}
\begin{defn}
Let $(U,E,\tau)$ be a IVHFSTS. Let $F_{A}$ be a IVHFSS over $(U,E).$
The interval-valued hesitant fuzzy soft closure of $F_A$ is defined as the intersection of
all interval-valued hesitant fuzzy soft closed sets(IVHFSCSs) which contained $F_A$ and is denoted by
$cl(F_A)$ or $\bar{F_A}.$ We write
\[\textrm{cl} (F_A)=\tilde{\bigcap}\{G_B: G_B\textrm {~is IVHFSCS
	and~} F_A\tilde{\subseteq}G_B\}.\]
\end{defn}
\begin{problem} 
From example \ref{prob35},  we have \\	$F_A=\{e_1=\{<h_1,[0.3, 0.8],[0.7,0.9]>, <h_2, [0.4, 0.6],[0.4,0.8],[0.5,0.7]>\},\\  e_2=\{<h_1,[0.6, 0.9], [0.7,1.0]>, <h_2, [0.3,0.6],[0.8,1.0]>\},\\
e_3=\{<h_1,[0.3, 0.6],[0.5,0.6] >, <h_2, [0.1,0.6],[0.3,0.9]>\}\}.$\\

$G_B=\{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8]>\},\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0] >, <h_2, [0.2,0.6],[0.8,1.0]>\}, \\e_3=\{<h_1,[0.0, 0.0],[0.0,0.0] >, <h_2, [0.0,0.0],[0.0,0.0]> \}\}.$\\
Then interval-valued hesitant fuzzy soft closed sets are\\
$F_A^C=\{e_1=\{<h_1,[0.1, 0.3],[0.2,0.7]>, <h_2, [0.3, 0.5],[0.2,0.6],[0.4,0.6]>\},\\  e_2=\{<h_1,[0.0, 0.3], [0.1,0.4]>, <h_2, [0.0,0.2],[0.4,0.7]>\},\\
e_3=\{<h_1,[0.4, 0.5],[0.4,0.7] >, <h_2, [0.1,0.7],[0.4,0.9]>\}\}.$\\

$G_B^C=\{e_1=\{<h_1,[0.2, 0.7]>, <h_2, [0.2, 0.7],[0.4,0.7]>\},\\  e_2=\{<h_1,[0.0, 0.3],[0.1,0.8] >, <h_2, [0.0,0.2],[0.4,0.8]>\},\\ e_3=\{<h_1,[1.0, 1.0],[1.0,1.0] >, <h_2, [1.0,1.0],[1.0,1.0]> \}\}.$\\
Suppose interval-valued hesitant fuzzy soft set $ I_C $ over $ (U,E) $ as\\
$I_C=\{e_1=\{<h_1,[0.1, 0.2],[0.1,0.7]>, <h_2, [0.3, 0.4],[0.1,0.6],[0.4,0.5]>\}\\  e_2=\{<h_1,[0.0, 0.2], [0.1,0.7]>, <h_2, [0.0,0.1],[0.4,0.8]>\}\\
e_3=\{<h_1,[0.0, 0.0],[0.0,0.0] >, <h_2, [0.0,0.0],[0.0,0.0]>\}\}.$\\

 Then\\
$cl(I_C)=\tilde{E} \tilde{\cap} G_B^C=G_B^C
=\{e_1=\{<h_1,[0.2, 0.7]>, <h_2, [0.2, 0.7],[0.4,0.7]>\},\\  e_2=\{<h_1,[0.0, 0.3],[0.1,0.8] >, <h_2, [0.0,0.2],[0.4,0.8]>\}, \\e_3=\{<h_1,[1.0, 1.0],[1.0,1.0] >, <h_2, [1.0,1.0],[1.0,1.0]> \}\}.$
\end{problem}
\begin{prop}
Let $ (U,E,\tau)$ be a IVHFSTS and $ F_A, G_B $ be two IVHFSs over $ (U,E).$ Then the following are true:
\begin{enumerate}
\item[(i)] $ \textrm{cl}(\tilde{\phi})=\tilde{\phi}, \textrm{cl}(\tilde{E})=\tilde{E} .$
\item[(ii)]$ F_A\tilde{\subseteq}\textrm{cl}(F_A) $
\item[(iii)] $ F_A $ is an interval-valued hesitant fuzzy soft closed set iff $ F_A= \textrm{cl}(F_A).$
\item[(iv)]$ F_A\tilde{\subseteq}G_B\Rightarrow\textrm{cl}(F_A)\tilde{\subseteq}\textrm{cl}(G_B)   $
\item[(v)]$\textrm{cl}(F_A\tilde{\cup}G_B)=\textrm{cl}(F_A)\tilde{\cup}\textrm{cl}(G_B)$
\item[(vi)]$\textrm{cl}(F_A\tilde{\cap}G_B)\tilde{\subseteq}\textrm{cl}(F_A)\tilde{\cap}\textrm{cl}(G_B)$	
\item[(vii)] $\textrm{cl}(\textrm{cl} F_A)= \textrm{cl}(F_A).$	
\end{enumerate}
\end{prop}
\begin{proof}
\begin{enumerate}
\item[(i)] Obvious. 
\item[(ii)] The proof directly follows from definition.
\item[(iii)] Let $ (U,E,\tau) $ be a IVHFSTS. Let $ F_A $ be a IVHFSS over $ (U,E) $ such that $ \textrm{cl}(F_A) =F_A. $\\
Therefore from definition of interval-valued hesitant fuzzy soft closure, we have $\textrm{cl}(F_A)  $ is interval-valued hesitant fuzzy soft closed sets. Hence $\textrm{cl}(F_A)  $ is interval-valued hesitant fuzzy soft closed and $ \textrm{cl}(F_A) =F_A. $ i.e $ F_A $ is  interval-valued hesitant fuzzy soft closed.\\
Conversely, let $ F_A $ be interval-valued hesitant fuzzy soft closed in $ (U,E,\tau). $ Therefore from definition of interval-valued hesitant fuzzy soft closure that any interval-valued hesitant fuzzy soft closed set $ G_B,  F_A\tilde{\subseteq}G_B\Rightarrow\textrm{cl}(F_A)\tilde{\subseteq}G_B . $\\
Since $ F_A\tilde{\subseteq}F_A\Rightarrow\textrm{cl}(F_A)\tilde{\subseteq}F_A  $ and from definition $F_A\tilde{\subseteq} \textrm{cl}(F_A) $\\
Hence it follows that $ F_A= \textrm{cl}(F_A).$
\item[(iv)] Let $ F_A\tilde{\subseteq}G_B. $  Since $ G_B\tilde{\subseteq}\textrm{cl}(G_B) .$ Therefore  $ F_A\tilde{\subseteq}\textrm{cl}(G_B) .$\\
Again $\textrm{cl}(F_A)  $ is the smallest interval-valued hesitant fuzzy soft closed set containing $ F_A. $\\
Hence $\textrm{cl}(F_A)\tilde{\subseteq}\textrm{cl}(G_B).$

\item[(v)] From definition of union of IVHFSSs\\
$ F_A\tilde{\subseteq}F_A\tilde{\cup}G_B, G_B\tilde{\subseteq}F_A\tilde{\cup}G_B. $\\
Therefore $ \textrm{cl}(F_A)\tilde{\subseteq}\textrm{cl}(F_A\tilde{\cup}G_B), \textrm{cl}(G_B)\tilde{\subseteq}\textrm{cl}(F_A\tilde{\cup}G_B)\\
\Rightarrow\textrm{cl}(F_A)\tilde{\cup}\textrm{cl}(G_B)\tilde{\subseteq}\textrm{cl}(F_A\tilde{\cup}G_B) $.......................................(A1)\\
Again $ \textrm{cl}(F_A\tilde{\cup}G_B)\tilde{\subseteq}\textrm{cl}(F_A)\tilde{\cup}\textrm{cl}(G_B) $.................................(A2)\\
Since $\textrm{cl}(F_A\tilde{\cup}G_B)  $ is the smallest interval-valued hesitant fuzzy soft closed set containing $F_A\tilde{\cup}G_B.  $
Hence from (A1) and (A2) , \\
$\textrm{cl}(F_A\tilde{\cup}G_B)=\textrm{cl}(F_A)\tilde{\cup}\textrm{cl}(G_B).$
\item[(vi)]From definition of intersection of IVHFSSs\\
$F_A\tilde{\cap}G_B\tilde{\subseteq} F_A, F_A\tilde{\cap}G_B\tilde{\subseteq}G_B.   $\\
Therefore\\ $\textrm{cl}(F_A\tilde{\cap}G_B)\tilde{\subseteq}\textrm{cl}(F_A), \textrm{cl}(F_A\tilde{\cap}G_B)\tilde{\subseteq}\textrm{cl}(G_B)\\
\Rightarrow \textrm{cl}(F_A\tilde{\cap}G_B)\tilde{\subseteq}\textrm{cl}(F_A)\tilde{\cap}\textrm{cl}(G_B). $
\item[(vii)] If $ F_A $ is a interval-valued hesitant fuzzy soft closed set then $  \textrm{cl}(F_A)=F_A .$ Hence 
$\textrm{cl}(\textrm{cl} F_A)= \textrm{cl}(F_A).$
\end{enumerate}
\end{proof}

\begin{defn} Let $(U,E,\tau)$ be a IVHFSTS. Let $F_{A}$ be a IVHFSS over $(U,E).$
	The interval-valued hesitant fuzzy soft  interior of $F_A$ is defined as the union of all
interval-valued hesitant fuzzy soft  open sets (IVHFSOSs) which contained $F_A$ and is denoted by
	$int(F_A)$ or $F^{\circ}_A.$ We write
	\[
	\textrm{int}(F_A)=\tilde{\cup}\{G_B: G_B \textrm {~is IVHFSOS and }G_B\tilde{\subseteq}F_A\}.\]
\end{defn}

\begin{problem} 
From example 3.5, we consider a interval-valued hesitant fuzzy soft set $ I_C $ over $ (U,E) $ as\\
$I_C=\{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.7],[0.3,0.8]>\}\\  e_2=\{<h_1,[0.2, 1.0], [0.7,1.0]>, <h_2, [0.2,0.7],[0.8,1.0]>\}\\
e_3=\{<h_1,[0.0, 0.0],[0.0,0.0] >, <h_2, [0.0,0.0],[0.0,0.0]>\}\}.$ \\
Therefore \\
$\textrm{int}(I_C)=G_B\tilde{\cup}\tilde{\phi}=G_B
= \{e_1=\{<h_1,[0.3, 0.8]>, <h_2, [0.3, 0.6],[0.3,0.8]>\},\\  e_2=\{<h_1,[0.2, 0.9],[0.7,1.0] >, <h_2, [0.2,0.6],[0.8,1.0]>,\\ e_3=\{<h_1,[0.0, 0.0],[0.0,0.0] >, <h_2, [0.0,0.0],[0.0,0.0]> \}\}.$			
\end{problem}
\begin{prop}
	Let $ (U,E,\tau)$ be a IVHFSTS and $ F_A, G_B $ be two IVHFSs over $ (U,E) .$ Then the following are true:
	\begin{enumerate}
		\item[(i)] $ \textrm{int}(\tilde{\phi})=\tilde{\phi}, \textrm{int}(\tilde{E})=\tilde{E} .$
		\item[(ii)]$\textrm{int}(F_A)\tilde{\subseteq} F_A $
		\item[(iii)] $ F_A $ is an interval-valued hesitant fuzzy soft open set iff $ F_A= \textrm{int}(F_A).$
		\item[(iv)]$ F_A\tilde{\subseteq}G_B\Rightarrow\textrm{int}(F_A)\tilde{\subseteq}\textrm{int}(G_B)   $
		\item[(v)]$\textrm{int}(F_A)\tilde{\cup}\textrm{int}(G_B)\tilde{\subseteq}\textrm{int}(F_A\tilde{\cup}G_B)$
		\item[(vi)]$\textrm{int}(F_A\tilde{\cap}G_B)=\textrm{int}(F_A)\tilde{\cap}\textrm{int}(G_B)$	
		\item[(vii)] $\textrm{int}(\textrm{int} F_A)= \textrm{int}(F_A).$	
	\end{enumerate}
\end{prop}
\begin{proof}
\begin{enumerate}
\item[(i)]	Obvious.
\item[(ii)] The proof directly follows from definition.
\item[(iii)] Let $ (U,E,\tau) $ be a IVHFSTS. Let $ F_A $ be a IVHFSS over $ (U,E) $ such that $ \textrm{int}(F_A) =F_A. $\\
Therefore from definition of interval-valued hesitant fuzzy soft interior, we have $\textrm{int}(F_A)  $ is interval-valued hesitant fuzzy soft open sets. Hence $\textrm{int}(F_A)  $ is interval-valued hesitant fuzzy soft open and $ \textrm{int}(F_A) =F_A. $ i.e $ F_A $ is  interval-valued hesitant fuzzy soft open.\\
Conversely, let $ F_A $ be an interval-valued hesitant fuzzy soft open in $ (U,E,\tau). $ Therefore from definition of interval-valued hesitant fuzzy soft interior that any interval-valued hesitant fuzzy soft open set $ G_B\tilde{\subseteq}F_A\Rightarrow G_B\tilde{\subseteq}\textrm{int}(F_A) . $\\
Since $ F_A\tilde{\subseteq}F_A\Rightarrow F_A\tilde{\subseteq}\textrm{int}(F_A)  $ and from definition $\textrm{int}(F_A)\tilde{\subseteq} F_A $\\
Hence, it follows that $ F_A= \textrm{int}(F_A).$
\item[(iv)] Let $ F_A\tilde{\subseteq}G_B. $ Since $\textrm{int}(F_A)\tilde{\subseteq}F_A \tilde{\subseteq}G_B, \textrm{int}(F_A)$ be a interval valued hesitant fuzzy soft open subset of $ G_B.$ Hence from definition of interval valued 
hesitant fuzzy soft interior, we have $ F_A\tilde{\subseteq}G_B\Rightarrow\textrm{int}(F_A)\tilde{\subseteq}\textrm{int}(G_B) .$
\item[(v)] Since $ F_A\tilde{\subseteq}F_A\tilde{\cup}G_B $ and 
 $ G_B\tilde{\subseteq}F_A\tilde{\cup}G_B .$ Therefore we have \\
 $ \textrm{int}(F_A)\tilde{\subseteq}\textrm{int}(F_A\tilde{\cup}G_B) $ and  $ \textrm{int}(G_B)\tilde{\subseteq}\textrm{int}(F_A\tilde{\cup}G_B).$ \\
 Hence
 $\textrm{int}(F_A)\tilde{\cup}\textrm{int}(G_B)\tilde{\subseteq}\textrm{int}(F_A\tilde{\cup}G_B).$
\item[(vi)] Since $F_A\tilde{\cap}G_B\tilde{\subseteq}F_A  $ and $ F_A\tilde{\cap}G_B\tilde{\subseteq}G_B. $ These implies that $ \textrm{int}(F_A\tilde{\cap}G_B)\tilde{\subseteq}\textrm{int}(F_A) $ and $ \textrm{int}(F_A\tilde{\cap}G_B)\tilde{\subseteq}\textrm{int}(G_B) . $\\
Therefore $\textrm{int}(F_A\tilde{\cap}G_B)\tilde{\subseteq} \textrm{int}(F_A)\tilde{\cap}\textrm{int}(G_B) . $......................(B1)\\
Again we know that $ \textrm{int}(F_A)\tilde{\subseteq}F_A $  and $ \textrm{int}(G_B)\tilde{\subseteq}G_B.  $ \\
Therefore $\textrm{int}(F_A)\tilde{\cap} \textrm{int}(G_B)\tilde{\subseteq}F_A\tilde{\cap}G_B..............................(B2) $\\
Hence from (B1) and (B2) we get \\ $\textrm{int}(F_A\tilde{\cap}G_B)=\textrm{int}(F_A)\tilde{\cap}\textrm{int}(G_B).$
\item[(vii)] From (iii), if $ F_A $ is an interval-valued hesitant fuzzy soft open set then $ F_A= \textrm{int}(F_A).$ Therefore 
$\textrm{int}(\textrm{int} F_A)= \textrm{int}(F_A).$	
\end{enumerate}
\end{proof}
\begin{prop}
If $ \{\tau_\lambda: \lambda\tilde{\in}I\} $ is a family of IVHFSTS on $ (U,E) ,$ then $\bigcap_\lambda \{\tau_\lambda: \lambda\tilde{\in}I\} $ is also a IVHFST on $ (U,E). $
\end{prop}
\begin{proof}
Suppose $ \{\tau_\lambda: \lambda\tilde{\in}I\} $  be a IVHFSTS. Therefore $ \tilde{\phi}, \tilde{E} \tilde{\in}\bigcap_{\lambda\tilde{\in}I} \{\tau_\lambda \}. $ If $ F_A ,G_B \tilde{\in}\bigcap_{\lambda\tilde{\in}I} \{\tau_\lambda \}$ then $ F_A, G_B \tilde{\in}\tau_\lambda, \forall \lambda \tilde{\in}I. $\\
Therefore $ F_A\tilde{\cap}G_B \tilde{\in}\tau_\lambda, \forall \lambda \tilde{\in}I. $\\
Thus $F_A\tilde{\cap}G_B \tilde{\in} \bigcap_{\lambda\tilde{\in}I} \{\tau_\lambda \}. $\\
Let $\{F_\alpha\}_{\alpha\tilde{\in}J}\tilde{\subseteq}\bigcap_{\lambda\tilde{\in}I} \{\tau_\lambda \}.$\\
Therefore $ F_\alpha\tilde{\subseteq}\bigcap_{\lambda\tilde{\in}I} \{\tau_\lambda \} ,\alpha\tilde{\in}J.$ This implies $F_\alpha\tilde{\subseteq}\tau_\lambda, \forall \lambda \tilde{\in}I, \alpha\tilde{\in}J.  $\\
Therefore $ \bigcup_{\alpha\tilde{\in}J}F_\alpha\tilde{\in}\bigcap_{\lambda\tilde{\in}I} \{\tau_\lambda \} $
\end{proof}
\begin{defn}
Let $ \tau_1 $ and $ \tau_2 $ be IVHFSTS on $ (U,E) .$ We say that $ \tau_1 $ is coarser (or weaker) than $ \tau_2 $ or $ \tau_2 $ is finer (or stronger) than $ \tau_1 $ if and only if $ \tau_1 \tilde{\subseteq}\tau_2 $ i.e. every $ \tau_1 $ interval valued hesitant fuzzy soft open set (IVHFSOS) is $ \tau_2 $ IVHFSOS. If either $ \tau_1 \tilde{\subseteq}\tau_2 $ or $ \tau_2 \tilde{\subseteq}\tau_1, $ we say that the IVHFST $ \tau_1 $ and  $ \tau_2 $ are comparable. 
If $\tau_1 \tilde{\nsubseteq}\tau_2   $ and $\tau_2 \tilde{\nsubseteq}\tau_1, $ we say the IVHFST $ \tau_1 $ and  
 $ \tau_2 $ are not comparable.
\end{defn}
\begin{problem}
From example 3.5, we consider IVHFST $ \tau_1 $ and $ \tau_2 $ on $ (U,E) $ as \\
$\tau_1=\{\tilde{\phi}, \tilde{E}, \tilde{F_A}\}.  $\\
$\tau_2=\{\tilde{\phi}, \tilde{E}, \tilde{F_A}, \tilde{G_B}\}.  $\\
Therefore $ \tau_1 \tilde{\subseteq}\tau_2 $ and hence $ \tau_1 $ is coarser than $ \tau_2. $
\end{problem}
\begin{defn}
The IVHFSS $ F_A $ over $ (U,E) $ is called a interval valued hesitant fuzzy soft point (IVHFSP) in $ (U,E) $ is denoted by $ e(F_A), $ if for the element $e\tilde{\in}A, \mu_{F(e)}\neq [0,0] $ and $ \mu_{F(e^{'})}= [0,0], \forall e^{'}\tilde{\in}A-{e}.$
\end{defn}
\begin{problem}\label{exa319}

Let $ U=\{h_1, h_2\} ,$ $ E=\{e_1, e_2, e_3\} $ and $ A=\{e_1, e_2\} \tilde{\subseteq}E. $ Suppose a IVHFSS $ F_A $ over $ (U,E) $ as\\
$F_A=\{e_1=\{<h_1,[0.0, 0.0]>, <h_2, [0.0, 0.0],[0.0,0.0]>\}\\  e_2=\{<h_1,[0.3, 0.5],[0.4,0.6] >, <h_2, [0.0,0.6]>\}\}.$\\
Here $ e_2\tilde{\in}A, \mu_{F(e_2)}\neq [0,0]  $ and for $ e^{'}\tilde{\in}A-{e_2}, \mu_{F(e^{'})}= [0,0]. $\\
Thus $ F_A $ is a IVHFSP in $ (U,E) $ denoted by $e_2(F_A).  $
\end{problem}
\begin{defn}
The IVHFSP $ e(F_A) $ is said to be in the IVHFSS $ G_B $ if $ A\tilde{\subseteq}B $ and for the element $ e\tilde{\in}A,\mu_{F(e)}\tilde{\subseteq}\mu_{G(e)}.  $ We denoted as $ e(F_A)\tilde{\in} G_B.$
\end{defn}
\begin{problem}\label{exam321}
From example \ref{exa319}, consider the IVHFSP $ e_2(F_A) $ and an IVHFSS $ G_B $ as \\
$G_B=\{e_1=\{<h_1,[0.1,0.9],[0.2, 0.3]>, <h_2, [0.6, 0.9]>\}\\  e_2=\{<h_1,[0.3, 0.6], [0.5,0.6],[0.5,0.8]>, <h_2, [0.3,0.8]>\}\\
e_3=\{<h_1,[0.5, 0.8] >, <h_2, [0.2,0.7],[0.3,0.8],[0.1,0.9]>\}\}.$\\
Here $ e_2\tilde{\in}A $ and $ \mu_{F(e_2)}\tilde{\subseteq}\mu_{G(e_2)}. $\\
Hence $  e_2(F_A) \tilde{\in}G_B.$
\end{problem}
\begin{defn}
A IVHFSS $ I_C $ in a IVHFSTS $ (U,E, \tau) $ is called a interval valued hesitant fuzzy soft neighborhood (IVHFSNBD) of the IVHFSP $ e(F_A)\tilde{\in}(U,E) $ if there is a IVHFSOS $ G_B $ such that $e(F_A)\tilde{\in}G_B\tilde{\subseteq}I_C.  $
\end{defn}
\begin{problem}\label{exa323}
From examples \ref{exa319}, \ref{exam321},  we consider the IVHFST $\tau=\{\tilde{\phi}, \tilde{E}, \tilde{G_B}\}.  $ and IVHFSS $ I_C $ as \\
$I_C=\{e_1=\{<h_1,[0.2,1.0]>, <h_2, [0.6, 0.9],[0.6,1.0]>\}\\  e_2=\{<h_1,[0.3, 0.6], [0.5,0.8]>, <h_2, [0.3,0.9]>\}\\
e_3=\{<h_1,[0.5, 0.8],[0.5,0.9] >, <h_2, [0.3,0.9],[0.3,1.0]>\}\\  e_4=\{<h_1,[0.1, 0.6], [0.7,0.9]>, <h_2, [0.2,0.6]>\}\},$\\ where $ E=\{e_1, e_2, e_3, e_4\},  C=\{e_1, e_2, e_3, e_4\} \tilde{\subseteq}E.$\\
Therefore  $e(F_A)\tilde{\in}G_B\tilde{\subseteq}I_C.  $\\
Hence $ I_C $ is a IVHFSNHD of the IVHFSP $e_2(F_A). $
\end{problem}
\begin{defn}
The family consisting of all neighborhoods of $ e(F_A)\tilde{\in}(U,E) $ neighborhood system of a fuzzy soft point $e(F_A). $ It is denoted by $ N_\tau(e(F_A)).$
\end{defn}
\begin{defn}
A IVHFSS $ I_C $ in a IVHFSTS $ (U,E,\tau) $ is called a IVHFSNBD of the IVHFSS $ H_A $ if there is a IVHFSOS $ G_B $ such that $ H_A \tilde{\subseteq}G_B\tilde{\subseteq}I_C. $
\end{defn}
\begin{problem}
From examples \ref{exam321}, \ref{exa323} and consider the IVHFSS $ H_A $ as \\
$H_A=\{e_1=\{<h_1,[0.1,0.5]>, <h_2, [0.6, 0.7],[0.6,0.8]>\}\\
e_3=\{<h_1,[0.5, 0.6],[0.4,0.6] >, <h_2, [0.2,0.3]>\}\},$\\ where $ A=\{e_1, e_3\} \tilde{\subseteq}E .$
Therefore $ H_A \tilde{\subseteq}G_B\tilde{\subseteq}I_C. $\\
Hence IVHFSS $ I_C $ is IVHFSNBD of the IVHFSS $ H_A. $
\end{problem}
\begin{prop}
The neighborhood system $ N_\tau(e(F_A))$ at, $ \forall e(F_A) $ in an IVHFSTS $ (U,E,\tau) $ has the following properties:
\begin{enumerate}
\item[(i)] If $ G_B \tilde{\in} N_\tau(e(F_A)) $ then $e(F_A)\tilde{\in}G_B.  $
\item[(ii)]If $ G_B \tilde{\in} N_\tau(e(F_A)) $ and $ G_B \tilde{\subseteq}H_C $ then $ H_C  \tilde{\in} N_\tau(e(F_A)).$
\item[(iii)]If $ G_B, H_C \tilde{\in} N_\tau(e(F_A)) $ then $ G_B\tilde{\cap} H_C \tilde{\in} N_\tau(e(F_A)) .$
\item[(iv)] If $ G_B \tilde{\in} N_\tau(e(F_A)) $ then there is a $H_C \tilde{\in} N_\tau(e(F_A))   $ such that  $ G_B \tilde{\in} N_\tau(e'(M_D)) $ for each $ e'(M_D)\tilde{\in}H_C. $
\end{enumerate}
\end{prop}
\begin{proof}
\begin{enumerate}
\item[(i)] If $ G_B \tilde{\in} N_\tau(e(F_A)), $ then there is a IVHFSOS $ H_C $ such that $ e(F_A)\tilde{\in}H_C\tilde{\subseteq}G_B. $ Therefore we have $e(F_A)\tilde{\in}G_B.  $
\item[(ii)] Let $ G_B \tilde{\in} N_\tau(e(F_A)), $ and $ G_B \tilde{\subseteq}H_C .$ Then there is a $ L_D $ such that $ e(F_A)\tilde{\in}L_D\tilde{\subseteq}G_B $ and $ e(F_A)\tilde{\in}L_D\tilde{\subseteq}G_B\tilde{\subseteq}H_C. $
Therefore $ H_C  \tilde{\in} N_\tau(e(F_A)).$
\item[(iii)] If $ G_B, H_C \tilde{\in} N_\tau(e(F_A)) $ then there exist IVHFSOSs $ L_D, M_E $ such that $ e(F_A)\tilde{\in}L_D\tilde{\subseteq}G_B $ and $ e(F_A)\tilde{\in}M_E\tilde{\subseteq}H_C. $ Thus $ e(F_A)\tilde{\in}L_D \tilde{\cap}M_E\tilde{\subseteq}G_B\tilde{\cap}H_C. $ Since $ L_D \tilde{\cap}M_E \tilde{\in}\tau. $ Hence we have $ G_B\tilde{\cap} H_C \tilde{\in} N_\tau(e(F_A)) .$ 
\item[(iv)] If $ G_B \tilde{\in} N_\tau(e(F_A)), $ then there is an IVHFSOS $ L_P \tilde{\in}\tau $ such that  $ e(F_A)\tilde{\in}L_P\tilde{\subseteq}G_B. $ Now put $ H_C=L_P. $ Then for each  $ e'(M_D)\tilde{\in}H_C,  e'(M_D)\tilde{\in}H_C\tilde{\subseteq}G_B. $ This implies 
 $ G_B \tilde{\in} N_\tau(e'(M_D)) .$ 
\end{enumerate}
\end{proof}

\end{document}